OpenStudy (anonymous):

Given: ΔBCD ~ ΔEFG Find x. A)4 B)8 C)10 D)12

OpenStudy (eyust707):

huh???

OpenStudy (anonymous):

here's the problem

OpenStudy (eyust707):

do you want me to teach u how to do it?

OpenStudy (anonymous):

yes please

OpenStudy (eyust707):

ok since the two triangles are similar if we can find the values of side that correspond to each other we can solve this...

OpenStudy (eyust707):

what is the length of CD

OpenStudy (anonymous):

idk,, do I add length CD or multiply?

OpenStudy (eyust707):

you have to use the pythagorean thrm

OpenStudy (eyust707):

a^2 + b^2 = c^2

OpenStudy (eyust707):

a and b are the legs... c is the hypontenuse

OpenStudy (eyust707):

for that triangle we know a and we know c... we need to solve for b

OpenStudy (eyust707):

still there?

OpenStudy (anonymous):

yes,, thinking

OpenStudy (eyust707):

kk

OpenStudy (anonymous):

a^2 + b^2 = c^2 6^2+? = ? ? + b^2 = Take from both sides: _______? b^2 = ? b = √? b = ? I really don't know

OpenStudy (eyust707):

yes very close

OpenStudy (eyust707):



OpenStudy (anonymous):

lol, can you explain

OpenStudy (eyust707):

yea definately

OpenStudy (eyust707):

do u have skype??

OpenStudy (anonymous):

no

OpenStudy (eyust707):

ohh ok well ill do my best on here

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

ok so if we have a right triangle and two sides we can use the pythagorean thrm to find the length of the unknown side